For given medium polarizing angle is 60 ,find critical angle

Grade:12th pass

2 Answers

Arun
6 years ago
tani = 1/sinc where i is the polarising angle and c is the critical angle
hence 
Sin c = 1/\sqrt3
C = sin -1(1/\sqrt3)
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
tani = 1/sinc where i is the polarising angle and c is the critical angle
hence 
Sin c = 1/√3
C = sin -1(1/√3)
 
Thanks and regards,
Kushagra